1. A full bathtub will empty at an average rate of 10 minutes. It will fill up to full at an average rate of 8 minutes if the plug is in. How long will it take to fill up to full if the plug is removed and the tap is turned on?

Answer by mananth(16946)

A full bathtub will empty at an average rate of 10 minutes. It will fill up to full at an average rate of 8 minutes if the plug is in. How long will it take to fill up to full if the plug is removed and the tap is turned on?
...
filling 8 minutes
1/8 of the job in 1 minute.
...
emptying 10 minutes
1/10 of this job in 1 minute.
..
1/8 -1/10 = effective filling in 1 minute
LCD = 40
(5-4)/40=1/40 of the job of filling in 1 minute.
So when plug is open it will take 40/1 minutes
40 minutes to fill the tub.
...
m.ananth@hotmail.ca